OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2 x + sin x = 0

OpenStudy (anonymous):

\[\sin^2x+\sin x=0\\(\sin x)(\sin x+1)=0\\\text{so either }\sin x\text{ is 0 or -1.}\\\sin x=0\\x=0,\pi\\\sin x=-1\\x=\frac{3\pi}2\]

OpenStudy (anonymous):

Thanks so much dude!

OpenStudy (anonymous):

No problem... do you see what I did? Factor the expression and set factors equal to 0 -- solve for \(x\).

OpenStudy (anonymous):

yeah, I see it now. Just couldn't remember it to save my life!